Suppose that you are pushing a child on a swing, initially at rest, periodically. Draw a displacement-time graph for the resulting oscillatory motion of the swing
1
Expert's answer
2020-02-27T10:03:27-0500

As we know that,

The swing will work as the simple pendulum, so as per the below diagram



Let the maximum displacement = a

So, here A is the amplitude,

let the time period of the swing is T,

"\phi=" phase difference

so "y= a\sin(\dfrac{2\pi}{T}t+\phi)"